OpenStudy (anonymous):

Simplify. Assume the divisor is not zero. (4x2y2 – 4xy) ÷ 4xy A. 2xy B. 2xy – y C. xy – 1 D. 2xy – 1

OpenStudy (anonymous):

@200205650 DDDDDDDDD

OpenStudy (anonymous):

No...you're close!

OpenStudy (anonymous):

BBBBBBBBBBBBB

OpenStudy (anonymous):

Lol no...

OpenStudy (anonymous):

\frac{ 8x^2y^2 -4xy}{ 4xy } \frac{ 8x^2y^2 }{ 4xy }-\frac{ 4xy }{ 4xy } 2x ^{2-1}y ^{2-1}-1 2xy – 1

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

DD

OpenStudy (anonymous):

No...you're close again!

OpenStudy (anonymous):

Literally!! XD

OpenStudy (anonymous):

really AAHHHHHHHHHH

OpenStudy (anonymous):

CCCCCCCCCCCCCCCC

OpenStudy (anonymous):

Yep! :)

OpenStudy (anonymous):

your really sure

OpenStudy (anonymous):

Ya y?

OpenStudy (anonymous):

*why

OpenStudy (anonymous):

nothing